OpenStudy (anonymous):

What is the slope of a line defined by the equation 3x + y = -4?

OpenStudy (anonymous):

Hi, cassb115! I think it's -3 as the slope.. :)

OpenStudy (anonymous):

Change the equation to slope intercept form to obtain the slope. 3x + y = -4 Subtract 3x from both sides. 3x + y = -4 -3x -3x y = -3x - 4 <=== In slope intercept form. ^ ^ Slope y - intercept So, the slope is -3. (:

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

No problem... ^-^